# A novel sensitivity analysis method for agent-based models stratifies in-silico tumor spheroid simulations

### v2

@@ 2.2.3 Reduce dimension
We use principal components analysis (PCA) to reduce the dimension of each DV, resulting in dimensionality-reduced DVs (DRDVs). We use scree plots to display the explained variance of each principal component and determine a suitable dimension for the DRDVs [37]. We use the python package Scikit-learn (version 1.4.2) to perform PCA.
@@ 2.2.4 Data clustering
We use $k$-means clustering [38], an unsupervised machine learning method, to group each DRDV into one of $k$ clusters. The $k$-means algorithm clusters $d$-dimensional data by identifying $k$ mean vectors, $\bar{x}_{1},\bar{x}_{2},\dots,\bar{x}_{k}\in\mathbb{R}^{d}$. Each data sample is assigned into the cluster whose mean it is closest to, using Euclidean distance. More information on how the mean vectors are found is provided in [38]. We use the elbow method to determine the number of clusters to use for each dataset [37]. We use the python package Scikit-learn (version 1.4.2) to perform clustering.
@@ 2.2.5 Statistical Analysis
Robustness assessment: We test the robustness of the clustering assignments by computing out-of-sample (OOS) accuracy scores. We achieve this by labeling each parameter value according to which cluster most of its samples are placed into. A high (low) OOS accuracy score suggests that many simulations from the same parameter value will be placed into the same cluster (different clusters).
@@
Computing OOS accuracy scores begins by splitting the 10 DRDVs from each parameter sample into 5 training and 5 testing DRDVs. We perform data clustering using the training DRDVs from all parameter values, and then label each parameter value according to the most common cluster assignments from its training DRDVs. To compute the OOS accuracy, we assign a “ground truth” label to each testing DRDV based on its parameter value’s label. We create a “predicted” label for each testing DRDV by running it through the trained clustering process. The OOS accuracy score is then given by the percentage of agreeing “ground truth” and “predicted” labels.
@@
Visual inspection: Each cluster from Section 2.2.4 contains DRDVs that were created over many parameters values. We use visualization to better understand which parameter values generate each cluster. When varying a small number of parameters (such as 2 or 3), we inspect a parameter partition plot, which colors each parameter value according to its cluster labeling [34]. When varying a larger number of parameters, we create ridgeline plots for each cluster that depict the distribution of parameter samples within each cluster. Each parameter distribution is estimated using kernel density estimation [39]. We use the python package Joypy (version 0.2.4) to create ridgeline plots.
@@
Statistical tests: For datasets generated by varying many parameters, we use the discrete two-sample Kolmogorov-Smirnov statistical test to evaluate whether the parameter distributions differ across clusters. For each varied parameter, we compare its samples between all (${k\choose 2}$ cluster pairs. Each comparison produces a $p$-value; low p-values suggest the parameter samples from the two clusters may be drawn from different distributions. We classify a parameter as impactful if the majority of these pairwise comparisons yield $p$-values below 0.05. We modified code from the ksdisc python package to perform statistical tests.
@@ 2.3 The Sobol’ GSA methodology
The Sobol’ Method is widely used to perform GSA [40, 41, 42]. We let $\bm{p}=(p_{1},p_{2},\dots,p_{Q})^{T}$ denote the input model parameters and $DV(\bm{p})$ represent the output DV summarizing an ABM simulation at the input parameter $\bm{p}$. The Sobol’ Method is a variance-based approach that estimates the contribution of each model parameter $p_{q}$ to the model’s total variance, $\text{Var}(DV(\bm{p}))$. Sobol’ introduced several sensitivity indices to quantify this contribution; we use the total-effect index in this study, which is given by [43, 44, 45]:
@@
where $\bm{p}_{\sim q}$ represents all parameters except for $p_{q}$, and $\text{E}_{\bm{p}_{\sim q}}(\cdot)$ and $\text{Var}_{\bm{p}_{\sim q}}(\cdot)$ represent the mean and variance over $\bm{p}$ while fixing $p_{q}$. This index quantifies the impact of changes to the input parameter $p_{q}$ on the model output $DV(\bm{p})$ while considering interactions between parameters.
The dimension of $S_{T_{q}}$ will match the dimension of $DV(\bm{p})$; suppose both are of length $m$. Our final sensitivity index is the averaged total effect index, which we calculate by averaging over all values in $S_{T_{q}}$:
$\bar{S}_{T_{q}}=\dfrac{1}{m}\sum_{i=1}^{m}(S_{T_{q}})_{i}.$ (2)
The values for $\bar{S}_{T_{q}}$ are typically estimated using quasi-Monte Carlo methods [46]. Saltelli [47, 48, 49] proposed an efficient sampling scheme to compute the total effect indices for all parameter values. This sampling requires $N(Q+2)$ parameter samples; it is common practice to choose $N$ to be a power of 2.
The Saltelli Dataset: As a comparison between the SSRCA and Sobol’ methods, we compute a third dataset generated using Saltelli parameter sampling [50]. Similar to Dataset 2, we generate this dataset by varying the same $Q=10$ parameters and choosing $N=2^{9}=512$ to ensure both datasets are of comparable size: Dataset 2 contains 5,445 samples while the Saltelli Dataset contains 6,144 samples. We implemented Saltelli sampling and the Sobol’ Index computations using the SALib package (version 1.5.1).
@@ 3 Results
3.1 Re-parameterizing the 2-dimensional Klowss Model
Computing the 3-dimensional ($\bm{x}=(x,y,z)$) Klowss Model across many parameter values is computationally prohibitive due to long simulation times. Instead, we simulate the 2-dimensional ($\bm{x}=(x,y)$) Klowss model, which corresponds to the equatorial cross section of the tumor spheroid at $z=0$. This reduction in dimension is justified by the radially symmetric nature of the model [21].
To account for this reduction, we modified two baseline parameter values: the number of starting agents, $N(0)$, and the ratio between nutrient consumption ($\kappa$) and diffusion ($D$) given by $\alpha=\kappa/D$. All other baseline values are set to those provided in [21]. To determine suitable $N(0)$ and $\alpha$ values, we simulated the 2-dimensional Klowss Model over a sample of both parameters and found that setting $N(0)=1100$ and $\alpha=0.01$ leads to close visual agreement with the $z=0$ cross section of the 3-dimensional Klowss Model (Supplementary Figures 7 and 8).
3.2 Summarizing the baseline model simulation
@@
Our baseline simulation refers to the model simulation computed using the baseline parameter values from Table 1. This simulation begins with all red, yellow, and green cells at time 0 (Figure 3). An outer ring of green cells forms by day 2, and a necrotic core has developed by day 4. The ring and core are clearly established by day 10, and the spheroid has grown in size over the 10-day simulation.
Figure 3: Descriptor vectors (DVs) of the baseline ABM simulation (a-d) Snapshots at times t= 0, 2, 6, and 10 days from the baseline simulation. e) The Cell counts DV for the baseline simulation. f) The Final cell density DV for the baseline simulation. The solid (dashed) circle has a radius of 137 $\mu m$ (322 $\mu m$).
@@
To summarize the baseline simulation, we compute its Cell counts and Final cell density DVs. According to the Cell counts DV, there is a steady increase in the number of red cells over the 10 day simulation, beginning with 850 and finishing with 3,000 cells, while the number of green and yellow cells remains relatively constant around 250 green cells and 10 yellow cells, respectively. The dead cell population initially grows slowly before increasing after $t=100$ hours; the simulation finishes with 1,000 dead cells. The Final cell density DV appears radially symmetric with an inner ring of 137 $\mu m$ and an outer ring of about 322 $\mu m$.
3.3 Analysis of Dataset 1
@@
We explore the two-dimensional Klowss Model’s behavior when varying two parameters, $\eta_{1}$ and $c_{a}$, to generate Dataset 1. We summarized all 1,000 simulations in this dataset by computing the Cell counts and Final cell density DVs. We then reduced the dimensionality of all DVs using PCA and determined that 3-dimensional DRDVs are suitable representatives for the DVs (Supplementary Fig S9) for both DV types.
@@
We trained the $k$-means clustering algorithm to the training datasets using both types of DRDVs separately and found that choosing $k=4$ or $5$ clusters can suitably cluster the data for both types of DRDVs. For simplicity, we present the results for $k=4$ clusters and note that the results are similar when setting $k=5$. When labeling each $(\eta_{1},c_{a})$ combination based on its most common clustering assignment, we find a clear partitioning of the $(\eta_{1},c_{a})$ parameter space into 4 distinct regions (Figure 4 and Supplementary Figures S11). The parameter partion plots that result from both types of DRDVs led to high OOS accuracy values, indicating that that these parameter designations are robust (Table 2).
@@
Figure 4: Dataset 1 Analysis using the Cell counts DV. (a) The ($\eta_{1},c_{a}$) parameter partition plot for Dataset 1 when using the Cell counts DV. Snapshots at times $t=2,6,\text{ and }10$ days from the representative model simulations from (b-d) Cluster 1, (e-g) Cluster 2, (h-j) Cluster 3, and (k-m) Cluster 4.
@@
Dataset 1 99.2% 98.0%
Dataset 2 99.6% 98.1%
Table 2: Out of sample (OOS) accuracy scores for both datasets when using either the Cell counts or Final cell density DRDVs.
Investigation of the parameter partition plot for the Cell counts DRDVs suggests that $c_{a}$ is the most sensitive parameter, as each cluster contains only some of the considered $c_{a}$ parameter values but all considered $\eta_{1}$ values (Figure 4). Cluster 1 results for $c_{a}\leq 0.232$, Cluster 2 results for $0.232\leq c_{a}\leq 0.348$, Cluster 3 results for $0.306\leq c_{a}\leq 0.606$, and Cluster 4 results for $c_{a}\geq 0.459$. We visualized representative model simulations from each cluster; the simulation from Cluster 1 results in the earliest-forming and largest final necrotic core, whereas the simulation from Cluster 4 has the smallest necrotic core, which emerges around 8 days. Visual inspection of the Cell counts DVs of these representative simulations reveal that as we increase the cluster assignment from 1 to 4, the number of red cells decreases and the number of dead cells decreases (Supplementary Figure S13). Similarly, the Final cell density DVs show that the ring of living cells decreases in radius as the clustering number increases.
We did not perform statistical analysis between the clusters from this dataset due to the ease in visually interpreting each parameter’s impact in Figure 4 and Supplementary Figure S11.
3.4 Analysis of Dataset 2
We further explore the two-dimensional Klowss Model’s behavior by sampling all 10 “Varying” parameters from Table 1 to generate Dataset 2. We summarized all 54,450 simulations in this dataset by computing the Cell count and Final cell density DVs and found that 3-dimensional DRDVs are suitable representatives for the DVs (Supplementary Figure 9), and that $k=4$ clusters reliably cluster the data (Supplementary Figure 10). We trained the $k$-means clustering algorithm to the training datasets for both types of DRDVs, and the parameter labels that resulted from both types of DRDVs led to high OOS accuracy values (Table 2). The representative simulations from the four clusters demonstrate that Cluster 1 (4) simulations have the largest (smallest) necrotic cores. Representative DVs for these simulations quantify these differences (Supplementary Figure 14).
Ridgeline plots visualize the distribution of parameter samples within each cluster (Figure 5). The ridgeline plots when using Cell counts or Final cell density DRDVs are similar, so we present the Cell counts DRDV results here (Final cell density DRDV results shown in Supplementary Figure S15). Visual inspection of the distributions reveals that:
@@ 3.4 Analysis of Dataset 2
  • Cluster 2 emerges from intermediate $c_{a}$ values, large $c_{d}$ values, low $\eta_{1}$ values, and low $\eta_{3}$ values;
  • Cluster 3 emerges from intermediate $c_{a}$ values, low to moderate $c_{d}$ values, and moderate to large $\eta_{1}$ values; and
@@
These observations suggest that the most sensitive models parameters are $c_{a}$, $c_{d}$, $\eta_{1}$, and $\eta_{3}$.
Figure 5: Dataset 2 Analysis using the Cell counts DV. (a-d) Ridgeline plots for Clusters 1-4. Each parameter $p$ is log-normalized as $\log_{2}(p/p_{base})$, where $p_{base}$ denotes the baseline model value from Table 1. Snapshots at times $t=2,6,\text{ and }10$ days from the representative model simulations from (e-g) Cluster 1, (h-j) Cluster 2, (k-m) Cluster 3, and (n-p) Cluster 4.
We identified sensitive model parameters by performing discrete two-sample Kolmogorov-Smirnov tests between all cluster pairs for all parameters when using the Cell counts DRDVs (Table 3). These tests confirm that $c_{a},c_{d},\eta_{1},\text{ and }\eta_{3}$ parameters are sensitive, and that all others parameters are insensitive. Performing the same sensitivity analysis using Final cell density DRDVs identifies these same four parameters are sensitive, as well as the $m_{min}$ parameter (Supplementary Table S5).
@@
$\eta_{3}$ $0.004$ $0.008$ 0.826 $<10^{-3}$ $<10^{-3}$ $<10^{-3}$ Sensitive
Table 3: Determining sensitive Dataset 2 parameters from the Cell counts DRDV. We designate $p$-values below 0.05 as being significant. If a parameter is significantly different for the majority of cluster pairs, then we classify the parameter as sensitive.
3.5 The Sobol’ and SSRCA Methods identify similar sensitive parameters
We compared the sensitivity analysis results that result from using the SSRCA and Sobol’ Methods. For the Sobol’ Method, we computed the averaged total effect effect indices, $\bar{S}_{T_{q}}$ from Equation (2), when varying all 10 parameters to generate the Saltelli Dataset. Both methods yield similar sensitivity analysis results. The $\bar{S}_{T_{q}}$ values are highest (indicative of high sensitivity) for the $c_{a}$ and $c_{d}$ parameters, followed by moderate $\bar{S}_{T_{q}}$ values for the $\eta_{1}$ and $\eta_{3}$ parameters; all other parameters have small $\bar{S}_{T_{q}}$ values (indicative of low sensitivity). The SSRCA method classifies these same four parameters as sensitive and all others as insensitive.
Figure 6: Comparing Sobol’ and SSRCA sensitivity analysis. We performed sensitivity analysis for the 10 model parameter using the Sobol’ indices methodology and the SSRCA methodology.

### v3

@@ 2.2.3 Reduce dimension
We use principal components analysis (PCA) to reduce the dimension of each DV, resulting in dimensionality-reduced DVs (DRDVs). We use scree plots to display the explained variance of each principal component and determine a suitable dimension for the DRDVs [james_introduction_2023]. We use the python package Scikit-learn (version 1.4.2) to perform PCA and data scaling.
Prior to performing PCA, we perform data scaling to standardize the DVs as follows. We do not perform data scaling for the Final cell density DV because different levels of variance are expected between different pixels. For example, some pixel values are zero for all simulations and have a variance of zero, while others will vary significantly between simulations and have large variance values. For the Cell counts DV, we first standardize each cell counts vector (e.g., $N_{r}(t),N_{y}(t),N_{g}(t),\text{ and }N_{d}(t)$) to have a variance of one (over all timepoints and data samples) before concatenating them together into the final DV. This is done to ensure that the PCA results are not corrupted by vectors with large magnitudes.
@@ 2.2.4 Data clustering
We use $k$-means clustering [kmeans], an unsupervised machine learning method, to group each DRDV into one of $k$ clusters. The $k$-means algorithm clusters $d$-dimensional data by identifying $k$ mean vectors, $\bar{x}_{1},\bar{x}_{2},\dots,\bar{x}_{k}\in\mathbb{R}^{d}$. Each data sample is assigned into the cluster whose mean it is closest to, using Euclidean distance. More information on how the mean vectors are found is provided in [kmeans]. We use the elbow method to determine the number of clusters to use for each dataset [james_introduction_2023]. We use the python package Scikit-learn (version 1.4.2) to perform clustering.
@@ 2.2.5 Statistical Analysis
Robustness assessment: We test the robustness of the clustering assignments by computing out-of-sample (OOS) consistency scores. We achieve this by labeling each parameter value according to which cluster most of its samples are placed into. A high (low) OOS consistency score suggests that many simulations from the same parameter value will be placed into the same cluster (different clusters).
@@
Computing OOS consistency scores begins by splitting the 10 DRDVs from each parameter sample into 5 training and 5 testing DRDVs. We perform data clustering using the training DRDVs from all parameter values, and then label each parameter value according to the most common cluster assignments from its training DRDVs. To compute the OOS consistency, we assign a “ground truth” label to each testing DRDV based on its parameter value’s label. We create a “predicted” label for each testing DRDV by running it through the trained clustering process. The OOS consistency score is then given by the percentage of agreeing “ground truth” and “predicted” labels. To avoid data leakage, the training and testing datasets are standardized separately, and only the training data is used to determine the principal components for PCA. The testing data is then projected onto the principal components determined from the training data.
@@
Visual inspection: Each cluster from Section 2.2.4 contains DRDVs that were created over many parameters values. We use visualization to better understand which parameter values generate each cluster. When varying a small number of parameters (such as 2 or 3), we inspect a parameter partition plot, which colors each parameter value according to its cluster labeling [angio_tda]. When varying a larger number of parameters, we create ridgeline plots for each cluster that depict the distribution of parameter samples within each cluster. Each parameter distribution is estimated using kernel density estimation [thrun_analyzing_2020]. To aid readers in interpreting ridgeline plots, we provide an illustrative example in Appendix S2 and Supplementary Figure S8. Briefly, sensitive parameters will have restricted domains and distinct distributions between clusters, whereas insensitive parameters will have wide domains and similar distributions between clusters. The heights of all distributions should be ignored in a ridgeline plot because all distributions are normalized to have a uniform height. We use the python package Joypy (version 0.2.4) to create ridgeline plots.
@@
Statistical tests: For datasets generated by varying many parameters, we use the discrete two-sample Kolmogorov-Smirnov statistical test to evaluate whether the parameter distributions differ across clusters. We choose the Kolomogorov-Smirnov test in place of other distribution-comparison tests (e.g., Mann-Whitney U test) due to its wide usage in biological studies and because it compares the entire distributions instead of only their mean or median locations. For each varied parameter, we compare its samples between all $k\choose 2$} cluster pairs. Each comparison produces a $p$-value; low p-values suggest the parameter samples from the two clusters may be drawn from different distributions. We classify a parameter as impactful if the majority of these pairwise comparisons yield $p$-values below 0.05. We modified code from the ksdisc python package to perform statistical tests.
2.3 Re-parameterizing the 2-dimensional Klowss Model
The Klowss Model from [4D_Tumor_Model] is 3-dimensional in space ($\bm{x}=(x,y,z)$), however, computing the model in all 3 spatial dimensions across many parameter values is computationally prohibitive due to long simulation times. We instead simulate the 2-dimensional ($\bm{x}=(x,y)$) Klowss model, corresponding to the equatorial cross section of the tumor spheroid at $z=0$. This reduction in dimension is justified by the radially symmetric nature of the model [4D_Tumor_Model].
To account for this reduction, we modified two baseline parameter values: the number of starting agents, $N(0)$, and the ratio between nutrient consumption ($\kappa$) and diffusion ($D$), given by $\alpha=\kappa/D$. All other baseline values are set to those provided in [4D_Tumor_Model], which we present in Table 1. We compared the resulting simulations of the 2-dimensional model to the $z=0$ cross section of the full 3-dimensional Klowss Model. We used the Cell counts descriptor vector (see Section 2.2.2) to summarize each simulation. From a small range of values for $N(0)$ and $\alpha$, we found that setting $N(0)=1100$ leads to good agreement on the number of cells in the G1 state for both models for early timepoints (Supplementary Figure S6), and setting $\alpha=0.01$ leads to good agreement between the two models for most cell states during later timepoints, although the 3-dimensional model contains more dead cells than the 2-dimensional model (Supplementary Figure S7).
@@ 2.4 The Sobol’ GSA Methodology
The Sobol’ Method is widely used to perform GSA [sobol_sensitivity_1990, sobol_sensitivity_1993, saltelli_relative_2002]. We let $\bm{p}=(p_{1},p_{2},\dots,p_{Q})^{T}$ denote the input model parameters and $DV(\bm{p})$ represent the output DV summarizing an ABM simulation at the input parameter $\bm{p}$. The Sobol’ Method is a variance-based approach that estimates the contribution of each model parameter $p_{q}$ to the model’s total variance, $\text{Var}(DV(\bm{p}))$. Sobol’ introduced several sensitivity indices to quantify this contribution. In particular, we consider the first-order and total-effect indices in this study, which are given by [smith_uncertainty_2013, saltelli_why_2019, saltelli_global_2007]:
$S_{q}=\dfrac{\text{Var}_{p_{q}}[\text{E}_{\bm{p}_{\sim q}}(DV(p_{q}))]}{\text{Var}(DV(\bm{p}))}\ \ q=1,2,\dots,Q,$ (1)
and
@@
respectively. Here, $\bm{p}_{\sim q}$ represents all parameters except for $p_{q}$, and $\text{E}_{\bm{p}_{\sim q}}(\cdot)$ and $\text{Var}_{\bm{p}_{\sim q}}(\cdot)$ represent the mean and variance over $\bm{p}$ while fixing $p_{q}$, respectively. Briefly, $S_{q}$ measures the direct contribution of the $p_{q}$ parameter to the model’s total variance, while $S_{T_{q}}$ measures the total contribution of the $p_{q}$ parameter to the variance both directly and indirectly through interactions with other parameters. Parameters with large $S_{q}$ and/or $S_{T_{q}}$ values can be considered sensitive and those with smaller values can be considered insensitive, although there are no clear guidelines on what constitutes a sensitive or insensitive parameter. Importantly, $S_{q}$ values are bound above by one, whereas $S_{T_{q}}$ values can exceed one because this index quantifies all interactions between 2 or more parameters.
The dimensions of $S_{q}$ and $S_{T_{q}}$ will match the dimension of $DV(\bm{p})$; suppose all are of length $m$. Our final sensitivity indices are then averaged first order and averaged total effect indices, which we calculate by averaging over all values in $S_{q}$ and $S_{T_{q}}$:
$\bar{S}_{q}=\dfrac{1}{m}\sum_{i=1}^{m}(S_{q})_{i},\ \text{ and }\ \ \bar{S}_{T_{q}}=\dfrac{1}{m}\sum_{i=1}^{m}(S_{T_{q}})_{i},\ \ \ q=1,2,\dots,Q$ (3)
The values for $\bar{S}_{q}$ and $\bar{S}_{T_{q}}$ are typically estimated using quasi-Monte Carlo methods [lima_bayesian_2021]. Saltelli [homma_importance_1996, saltelli_making_2002, saltelli_variance_2010] proposed an efficient sampling scheme to compute the total effect indices for all parameter values. This sampling requires $N(Q+2)$ parameter samples; it is common practice to choose $N$ to be a power of 2.
The Saltelli Dataset: As a comparison between the SSRCA and Sobol’ Methods, we compute a third dataset generated using Saltelli parameter sampling [saltelli_sensitivity_2002]. Similar to the large 10-parameter dataset, we generate this dataset by varying the same $Q=10$ parameters and choosing $N=2^{9}=512$ to ensure both datasets are of comparable size: the large 10-parameter dataset contains 5,445 samples while the Saltelli Dataset contains 6,144 samples. We implemented Saltelli sampling and the Sobol’ Index computations using the SALib package (version 1.5.1).
@@ 3 Results
3.1 Quantifying the baseline simulation’s proliferative ring and necrotic core
@@
Our baseline simulation refers to the model simulation computed using the baseline parameter values from Table 1. This simulation begins with all cells in the three living states at time 0 (Figure 2(c)). An outer ring of proliferating cells forms by day 2, and a necrotic core has developed by day 4. The ring and core are clearly established by day 10, and the spheroid has grown in size over the 10-day simulation.
@@
To summarize the baseline simulation, we compute its Cell counts and Final cell density DVs. According to the Cell counts DV, there is a steady increase in the number of cells in the G1 state over the 10 day simulation, beginning with 850 and finishing with 3,000 cells, while the number of proliferating cells remains relatively constant around 10 cells in the early S state and 250 cells in the S/G2/M state (Figure 2(d)). The dead cell population initially grows slowly before increasing after $t=4$ days; the simulation finishes with 1,000 dead cells. The Final cell density DV appears radially symmetric with an inner ring of 137 $\mu m$ and an outer ring of about 322 $\mu m$ (Figure 2(e)).
3.2 SSRCA identifies four model phenotypes within the simple 2-parameter dataset
@@
We explore the two-dimensional Klowss Model’s behavior when varying two parameters, $\eta_{1}$ and $c_{a}$, to generate the simple 2-parameter dataset. We summarized all 1,210 simulations in this dataset by computing the Cell counts and Final cell density DVs. We then reduced the dimensionality of all DVs using PCA and determined that 3-dimensional DRDVs are suitable representatives for the DVs (Supplementary Figure S9(a)-(b)) for both DV types.
@@
We trained the $k$-means clustering algorithm to the training datasets using both types of DRDVs separately. Elbow plots suggest there may be either $k=4$ or $k=5$ clusters in the data (Supplementary Figure S10(a)-(b)), we found that either choice leads to similar results. We proceed with the simpler $k=4$ scenario moving forward. When labeling each $(\eta_{1},c_{a})$ combination based on its most common clustering assignment, we find a clear partitioning of the $(\eta_{1},c_{a})$ parameter space into four distinct regions (Figure 3(a) and Supplementary Figure S11(a)). The parameter partition plots that result from both types of DRDVs led to high OOS consistency values, indicating that that these parameter designations are robust (Table 2).
@@
Figure 3: The simple 2-parameter dataset Analysis using the Cell counts DV. (a) The ($\eta_{1},c_{a}$) parameter partition plot for the simple 2-parameter dataset when using the Cell counts DV. Snapshots at times $t=2,6,\text{ and }10$ days from the representative model simulations from (b-d) Cluster 1, (e-g) Cluster 2, (h-j) Cluster 3, and (k-m) Cluster 4.
@@
The simple 2-parameter dataset 99.2% 98.0%
The large 10-parameter dataset 99.6% 98.1%
Table 2: Out of sample (OOS) consistency scores for both datasets when using either the Cell counts or Final cell density DRDVs.
Investigation of the parameter partition plot for the Cell counts DRDVs reveals that each cluster contains only a subset of the $c_{a}$ values (Figure 3(a)). In contrast, all clusters include the full range of $\eta_{1}$ values. These results suggest that $c_{a}$ impacts the model’s behavior more than $\eta_{1}$. Cluster 1 results for $c_{a}\leq 0.232$, Cluster 2 results for $0.232\leq c_{a}\leq 0.348$, Cluster 3 results for $0.306\leq c_{a}\leq 0.606$, and Cluster 4 results for $c_{a}\geq 0.459$. We visualized representative model simulations from each cluster (Figure 3(b)-(m)); the simulation from Cluster 1 results in the earliest-forming and largest final necrotic core, whereas the simulation from Cluster 4 has the smallest and last-forming necrotic core. Visual inspection of the Cell counts DVs of these representative simulations reveal that as we increase the cluster assignment from 1 to 4, the number of G1 cells decreases and the number of dead cells decreases (Supplementary Figure S12(a)-(d)). Similarly, the Final cell density DV leads to similar representative model simulations (Supplementary Figure S11(b)-(m)), and the representative DVs show that the ring of living cells decreases in radius as the clustering assignment increases (Supplementary Figure S12(e)-(h)).
We did not perform statistical analysis between the clusters from this dataset due to the ease in visually interpreting each parameter’s impact in Figure 3(a) and Supplementary Figure S11(a).
3.3 SSRCA identifies sensitive cell cycle- and death-related parameters from the large 10-parameter dataset
We further explore the two-dimensional Klowss Model’s behavior by sampling all 10 “Varying” parameters from Table 1 to generate the large 10-parameter dataset. We summarized all 54,450 simulations in this dataset by computing the Cell count and Final cell density DVs and found that 3-dimensional DRDVs are suitable representatives for the DVs (Supplementary Figure S9(c)-(d)), and that $k=4$ reliably clusters the data (Supplementary Figure S10(c)-(d)). We trained the $k$-means clustering algorithm to the training datasets for both types of DRDVs, and the parameter labels that resulted from both types of DRDVs led to high OOS consistency values (Table 2).
Ridgeline plots visualize the distribution of parameter samples for all four clusters. The ridgeline plots when using Cell counts (Figure 4(a)-(d)) or Final cell density DRDVs (Supplementary Figure S13(a)-(d)) are similar, so we present the Cell counts DRDV results here for simplicity. The distributions for the $c_{a}$, $c_{d}$, $\eta_{1}$, and $\eta_{3}$ parameters each have distinct domains between the four clusters, which indicates they are sensitive parameters. The remaining six parameters have similar domains for all four clusters and are likely insensitive. For the four sensitive parameters, we notice that:
@@
  • Cluster 2 emerges from low-to-moderate $c_{a}$ values, large $c_{d}$ values, low $\eta_{1}$ values, and low $\eta_{3}$ values;
  • Cluster 3 emerges from moderate-to-large $c_{a}$ values, low to moderate $c_{d}$ values, and moderate to large $\eta_{1}$ values; and
@@
We visualize representative simulations from each cluster (Figure 4(e)-(p)). The Cluster 1 simulation has the largest necrotic core and thinnest proliferating ring. The size of the core shrinks and the thickness of the ring increases for the simulations from Clusters 2 and 3. The Cluster 4 simulation does not include a necrotic core and consists primarily of living cells. The DVs for for each representative simulation quantify these differences (Supplementary Figure S14(a)-(d)). We find similar representative simulations when using the Final cell density DV (Supplementary Figure S13(e)-(p)), and their respective DVs show that the simulations’ radii decrease as the clustering assignment increases (Supplementary Figure S14(e)-(h)).
Figure 4: The large 10-parameter dataset Analysis using the Cell counts DV. (a-d) Ridgeline plots for Clusters 1-4. Each parameter $p$ is log-normalized as $\log_{2}(p/p_{base})$, where $p_{base}$ denotes the baseline model value from Table 1. Snapshots at times $t=2,6,\text{ and }10$ days from the representative model simulations from (e-g) Cluster 1, (h-j) Cluster 2, (k-m) Cluster 3, and (n-p) Cluster 4.
We confirmed the sensitive model parameters for this dataset by performing discrete two-sample Kolmogorov-Smirnov tests between all cluster pairs for all parameters when using the Cell counts DRDVs (Table 3). The tests find that the $c_{a},\ c_{d},\ \eta_{1},\text{ and }\eta_{3}$ parameters achieve distributions with low $p$-values between the majority of cluster pairs. This suggests that the distributions of these parameters are different between clusters, which we interpret as parameter sensitivity. All other parameters achieve higher $p$-values between cluster pairs, indicative of similar distributions between clusters and parameter insensitivity. Performing the same sensitivity analysis using Final cell density DRDVs identifies these same four parameters are sensitive, as well as the $m_{min}$ parameter (Supplementary Table S5).
@@
$\eta_{3}$ $0.004$ $0.008$ 0.826 $10^{-3} <$$10^{-3} <$$10^{-3}$ Sensitive
Table 3: Determining sensitive parameters from the large 10-parameter dataset when using the Cell counts DRDV. A small $p$-value between two distributions indicates they are significantly different from each other. We designate $p$-values below 0.05 as being significant. If a parameter is significantly different for the majority of cluster pairs, then we classify the parameter as sensitive.
3.4 SSRCA provides stable sensitivity analysis results while Sobol’ does not
[…]
Figure 5: Comparing Sobol’ and SSRCA sensitivity analysis. We performed sensitivity analysis for the 10 model parameter using the Sobol’ indices methodology and the SSRCA Methodology when using the (a) Cell counts DV and (b) the Final cell density DV.
There is less agreement between the two methods when using the Final cell density DV (Figure 5(b)). The Sobol’ Method leads to all ten parameters achieving $\bar{S}_{T_{q}}$ values of 0.58 or greater, which could suggest that all parameters are sensitive. However, inspection of the $\bar{S}_{q}$ values reveals that only the $c_{a}$ parameter achieves a sensitivity index value above 0.05, while all others are at 0.04 or below. The large discrepancy between the $\bar{S}_{q}$ and $\bar{S}_{T_{q}}$ index values for the nine other parameters suggests that the model variance is highly dependent on interactions between parameters. Using the SSRCA Methodology with the Final cell density DVs, on the other hand, designates five total sensitive parameters: $c_{a},c_{d},\eta_{3},\eta_{1}$, and $m_{\text{min}}$. Importantly, these are the same parameters that SSRCA designated for the Cell counts DV with the addition of the $m_{\text{min}}$ parameter. This fifth parameter may be a result of the Final cell density DV capturing a spatial aspect of the model that is missing from the Cell counts DV.